### v4

# Octic Vision Transformers: Quicker ViTs Through Equivariance

Abstract
Why are state-of-the-art Vision Transformers (ViTs) not designed to exploit natural geometric symmetries such as 90-degree rotations and reflections? In this paper, we argue that there is no fundamental reason, and what has been missing is an efficient implementation. To this end, we introduce Octic Vision Transformers (octic ViTs) which rely on octic group equivariance to capture these symmetries. In contrast to prior equivariant models that increase computational cost, our octic linear layers achieve 5.33x reductions in FLOPs and up to 8x reductions in memory compared to ordinary linear layers. In full octic ViT blocks the computational reductions approach the reductions in the linear layers with increased embedding dimension. We study two new families of ViTs, built from octic blocks, that are either fully octic equivariant or break equivariance in the last part of the network. Training octic ViTs supervised (DeiT-III) and unsupervised (DINOv2) on ImageNet-1K, we find that they match baseline accuracy while at the same time providing substantial efficiency gains.
1 Introduction
In the pursuit of flexible yet scalable models, Vision Transformers (ViTs) (Dosovitskiy et al., 2021) have emerged as the dominant architecture in modern computer vision. Key to their success is the combination of visual tokens, constructed from image patches, with the powerful attention mechanism (Vaswani et al., 2017), resulting in a versatile and scalable architecture. This scalability is due in large part to weight-sharing between the tokens, which ensures permutation equivariance.
Equivariance provides a powerful inductive bias in neural networks by enforcing structured responses to transformations such as permutations, translations, rotations or reflections. Another major benefit of equivariance is the potential for reducing computational costs, for instance by token-wise weight-sharing as mentioned, or by parameterizing networks in the Fourier domain of a symmetry group. It is therefore interesting to investigate whether imbuing ViTs with further equivariance can yield improved compute efficiency. Equivariance to roto-reflections, formalized through the octic group $\mathrm{D}_{8}$, is a particularly attractive group, as many vision tasks exhibit such symmetries.
$\mathrm{D}_{8}$ equivariance was introduced to Convolutional Neural Networks (CNNs) by Cohen & Welling (2016), demonstrating improved parameter efficiency through weight sharing (Wood & Shawe-Taylor, 1996; Bekkers et al., 2018; Weiler & Cesa, 2019). Yet, despite its theoretical appeal, state-of-the-art vision models (Radford et al., 2021; Kirillov et al., 2023; Oquab et al., 2024; Wang et al., 2024a) do not incorporate group equivariance other than the permutation equivariance of transformer layers. We argue that this is not due to a lack of utility, but due to practical limitations: most existing implementations construct equivariant layers using computationally inefficient architectures, leading to increased FLOPs and runtime. Without efficient, hardware-compatible implementations, these methods have remained impractical for large-scale models. As a result, the potential of equivariant design remains largely unexplored in state-of-the-art systems, in particular in the context of ViTs. Recently, Bökman et al. (2025) showed that equipping ViTs with horizontal flip equivariance results in retained performance while saving FLOPs. However, their improvements are limited due to the small cardinality of their chosen group.
In this paper, we demonstrate that scaling equivariance to larger groups can be efficiently implemented – yielding faster, stronger, and more compact models, cf. Figure 1. Specifically, we introduce octic-equivariant layers for ViTs, leveraging the $\mathrm{D}_{8}$ symmetry group of $90^{\circ}$ rotations and reflections. Our approach integrates seamlessly into existing ViT architectures and leads to significant gains in throughput and memory efficiency, without sacrificing accuracy.
In summary, our contributions are as follows:
  1. (a)
    We introduce octic-equivariant layers for ViTs, described in Section˜3. In our implementation, octic-equivariant linear layers use $5.33$ times fewer FLOPs and $8$ times less memory per feature dimension than ordinary linear layers. Octic-equivariant ViT layers hence asymptotically have the same compute savings (Table˜1, Figure˜4(a)).
  2. (b)
    We propose two highly competitive families of ViTs, that are either fully rotation equivariant ($\mathcal{I}_{8}$) or break equivariance in late layers of the network ($\mathcal{H}_{8}$).
  3. (c)
    In Section˜4, we demonstrate empirically that our ViTs can be used in state-of-the-art ViT training recipes (DeiT III and DINOv2) without re-tuning hyperparameters. In particular, we achieve a $40\%$ FLOP saving with our $\mathcal{I}_{8}(\text{ViT-H})$ and $\mathcal{H}_{8}(\text{ViT-H})$ models while matching baseline performance (Figure˜1).
  4. (d)
    We study the effects of different methods of invariantization (testing six different methods in Appendix˜D), completely breaking equivariance (Section˜4.3) and varying the number of equivariant layers (Section˜4.3). These ablations will help guide future research on equivariant architectures at scale.
In addition to introducing new state-of-the-art ViTs, from a broader perspective our contributions give clear evidence that equivariance can matter at scale, which has been a subject of debate in recent literature (Abramson et al., 2024; Wang et al., 2024b; Brehmer et al., 2025; Bökman et al., 2025).
Figure 1: Computational savings. Using octic layers in ViTs significantly reduces the computational complexity without sacrificing accuracy on ImageNet-1K, for both supervised and self-supervised training. Detailed results can be found in Section 4.
2 Related Work
Vision Transformers.
The ViT was introduced by Dosovitskiy et al. (2021) and has subsequently achieved state-of-the-art results in many domains of computer vision (Carion et al., 2020; Radford et al., 2021; Kirillov et al., 2023; Edstedt et al., 2024; Wang et al., 2025). Significant efforts have been made to scale ViTs (Zhai et al., 2022; Dehghani et al., 2023), alongside strategies to do so efficiently (Alabdulmohsin et al., 2023). In this work, we propose incorporating octic layers in ViTs, which not only maintain efficiency at larger scales but become increasingly effective as model size grows, thus directly leveraging the scaling of ViTs. Hierarchical Transformers (Liu et al., 2021; Wu et al., 2021; Hassani et al., 2023) use translational symmetry, and SparseViT (Chen et al., 2023) extends this with sparse activations. In contrast, our work instead focuses on roto-reflections and builds exact equivariance into the architecture.
Equivariant Networks.
The equivariance of CNNs to (cyclic) image translations can be extended to incorporate larger symmetry groups such as rotations and reflections, as shown by Cohen & Welling (2016); Dieleman et al. (2016) using Group Equivariant CNNs (G-CNNs). Cohen & Welling (2017); Weiler & Cesa (2019) generalized G-CNNs to steerable CNNs, where the features transform according to general group representations. Our octic ViTs can be seen as a more scalable ViT analogue of the octic steerable CNNs by Cohen & Welling (2017). There have also been prior efforts on attention- and Transformer-based equivariant architectures. For both point clouds (Fuchs et al., 2020; Hutchinson et al., 2021; Assaad et al., 2023; Liao & Smidt, 2023) and, more closely to ours, images (Romero et al., 2020; Xu et al., 2023; Rojas-Gomez et al., 2024; Kundu & Kondor, 2025). However, these prior works do not obtain computational benefits over non-equivariant transformers, in contrast to our ViTs. Our work is part of an ongoing research direction of studying and improving the scalability of equivariant networks (Bekkers et al., 2024; Brehmer et al., 2025; Bharadwaj et al., 2025; Vadgama et al., 2025; NVIDIA, ). Prior work in this direction mostly focused on point cloud data, with the notable exception of He et al. (2021) and Bökman et al. (2025) who considered images. He et al. (2021) increased the computational efficiency of G-CNNs but in contrast to our work did not achieve such benefits against standard networks. Directly inspiring our work, Bökman et al. (2025) demonstrated that incorporating horizontal mirroring equivariance into modern image classifiers increases compute efficiency while maintaining representational power. In contrast, we consider the larger octic group and thus achieve further savings in FLOPs. We conduct more extensive experimentation than Bökman et al. (2025) and address open questions in their work by studying the effect of breaking equivariance, invariantization and the number of equivariant blocks.
3 Method
In this section, we design octic-equivariant ViT layers. We begin with preliminaries for octic equivariance in Section 3.1, followed by the introduction of octic ViTs in Section 3.2, specifics of the Transformer layers in Section 3.3, and a detailed discussion of computational efficiency in Section 3.4. We summarize the most important notation in Section 3.1.1.
3.1 Preliminaries on Octic Equivariance
In this work we focus on the dihedral group with eight elements, $\mathrm{D}_{8}$ $\{e,r,r^{2},r^{3},s,sr,sr^{2},sr^{3}\}$, such that $r^{4}=s^{2}=e$ is the identity element and $r^{3}=srs$. [^1] The reader is cautioned that $\mathrm{D}_{8}$ is sometimes alternatively denoted $\mathrm{D}_{4}$. We think of $\mathrm{D}_{8}$ as acting on images by reflections $s$ and 90° rotations $r$. $\mathrm{D}_{8}$ is also called the octic group, and we opt for this shorter name throughout.
We consider network layers as maps between real group representations of $\mathrm{D}_{8}$. In our setting, a real representation is a vector space $\mathbb{R}^{n}$ equipped with a group homomorphism $\rho$ from $\mathrm{D}_{8}$ to the group of $n\times n$ invertible real matrices. In other words for every $g\in\mathrm{D}_{8}$, $\rho(g)$ is an invertible matrix and for every $g,h\in\mathrm{D}_{8}$, $\rho(gh)=\rho(g)\rho(h)$. $\rho$ specifies how $x\in\mathbb{R}^{n}$ transforms under $\mathrm{D}_{8}$. A representation of $\mathrm{D}_{8}$ is also defined by choosing matrices $\rho(r)$ and $\rho(s)$ such that $\rho(r)^{4}=\rho(s)^{2}=I$ and $\rho(r)^{-1}=\rho(s)\rho(r)\rho(s)$. There are only a few different representations of $\mathrm{D}_{8}$ used in this work, we list them below as examples. Finally, it is worth mentioning that all representations considered here are orthogonal, i.e., satisfy $\rho(g)^{-1}=\rho(g)^{\mathsf{T}}$.
The atomic building blocks of group representations are the so-called irreducible representations.
Example 3.1 (Irreducible representations).
The five irreducible representations, short irreps, of $\mathrm{D}_{8}$ are defined by
[…]
We use the same notation as the original work on steerable CNNs (Cohen & Welling, 2017) for these irreps, but choose a different basis for $\rho_{E}$. It is known from elementary representation theory that any representation of $\mathrm{D}_{8}$ can be decomposed into irreps as
$\rho(g)=Q\left(\raisebox{0.86108pt}{$\bigoplus_{i\in\{\mathrm{A1},\mathrm{A2},\mathrm{B1},\mathrm{B2},\mathrm{E}\}}m_{i}\rho_{i}(g)$}\right)Q^{-1}$ (2)
where $\oplus$ denotes direct sum of representations, or stacking matrices in a block diagonal, and we write $m_{i}\rho_{i}(g)$ for $\oplus$’ing $\rho_{i}(g)$ with itself $m_{i}$ times. Here $Q$ is an invertible matrix and the $m_{i}$ are integers specifying the multiplicity of each irrep.
Example 3.2 (Regular representation).
The regular representation $\rho_{\text{reg}}$ can be thought of as $\mathrm{D}_{8}$ acting canonically on the vector space of functions $\phi:\mathrm{D}_{8}\to\mathbb{R}$:
$\left[\rho_{\text{reg}}(g)\phi\right](h)=\phi(g^{-1}h).$ (3)
We identify each $\phi:\mathrm{D}_{8}\to\mathbb{R}$ with the vector
$\begin{pmatrix}\phi(e)&\phi(r^{3})&\phi(r^{2})&\phi(r)&\phi(s)&\phi(sr^{3})&\phi(sr^{2})&\phi(sr)\end{pmatrix}^{\mathsf{T}}\in\mathbb{R}^{8}$ (4)
so that $\rho_{\text{reg}}(g)$ can be written as a permutation matrix. Importantly, $\rho_{\text{reg}}$ commutes with pointwise activation functions such as GELU (Hendrycks & Gimpel, 2016).
Example 3.3 (Isotypical decomposition / Fourier transform).
The regular representation $\rho_{\text{reg}}$ can be block-diagonalized to its isotypical decomposition $\rho_{\text{iso}}$ through equation 2 as $\rho_{\text{reg}}(g)=Q_{\text{reg}}\rho_{\text{iso}}(g)Q_{\text{reg}}^{-1}$ with
$\rho_{\text{iso}}(g)=\rho_{\text{A1}}(g)\oplus\rho_{\text{A2}}(g)\oplus\rho_{\text{B1}}(g)\oplus\rho_{\text{B2}}(g)\oplus 2\rho_{\text{E}}(g).$ (5)
It is a general fact for finite groups that the regular representation decomposes into all irreps with multiplicities equal to their dimensions. The change of basis $Q_{\text{reg}}$ (written out in full in Appendix˜A) is the inverse Fourier transform of $\mathrm{D}_{8}$, with $Q_{\text{reg}}^{-1}=Q_{\text{reg}}^{\mathsf{T}}$ being the Fourier transform.
[…]
Figure 2: $\mathrm{D}_{8}$ Linear layers. Implementing equivariant linear layers in the Fourier domain of $\mathrm{D}_{8}$ gives a major computational benefit. Left: A $C\times C$ weight matrix being multiplied by $L$ tokens of feature dimension $C$. Center: The block-diagonalization that happens when enforcing the layer to be $\mathrm{D}_{8}$-equivariant in the Fourier domain. More precisely, we enforce equivariance with respect to the representation $\frac{C}{8}\rho_{\text{iso}}$ that splits into irreps $\rho_{\text{A1}},\rho_{\text{A2}},\rho_{\text{B1}},\rho_{\text{B2}}$ and $\rho_{\text{E}}$ as detailed in Section 3. There is no mixing between different irreps and the weight sharing in the block-diagonal stems from the fact that $\rho_{\text{E}}$ is a two-dimensional irrep. Right: An efficient implementation of the original $C\times C$ by $C\times L$ matrix multiplication as four $\frac{C}{8}\times\frac{C}{8}$ by $\frac{C}{8}\times L$ and one $\frac{C}{4}\times\frac{C}{4}$ by $\frac{C}{4}\times 2L$ matrix multiplication. An equivariant linear layer of this type requires $16/3\approx 5.33$ times fewer FLOPs to compute and has 88 times fewer parameters than the ordinary linear layer shown to the left.
Equivariant networks use equivariant linear layers, which map between representations. A linear $G$-equivariant map or intertwiner, $W\in\mathbb{R}^{n\times n}$, between representations $\rho_{1},\rho_{2}$, commutes with their action, i.e., $\rho_{2}(g)W=W\rho_{1}(g)$. It follows from Schur’s lemma that $W=\lambda I$ for some scalar $\lambda$ if $\rho_{1},\rho_{2}$ are irreps, and that $\lambda=0$ if $\rho_{1},\rho_{2}$ are not isomorphic (Serre, 1977, Section I.2.2) [^2] We can apply Schur’s lemma for complex representations here since the irreps listed in Example 3.1 are irreducible over the complex numbers. We will however only use real-valued linear layers, i.e. $\lambda\in\mathbb{R}$.. As $\rho_{\text{iso}}$ is just a stack of irreducible representations, any intertwiner between such representations will be sparse, which is why they are less computationally expensive than ordinary linear layers. The naive computational complexity for a linear map $W:\mathbb{R}^{|\mathrm{D}_{8}|}\to\mathbb{R}^{|\mathrm{D}_{8}|}$ is $|\mathrm{D}_{8}|^{2}=64$ multiplications [^3] We ignore the additions for simplicity.. In contrast, intertwiners $\rho_{\text{iso}}\to\rho_{\text{iso}}$ require a total of $\sum_{i}^{k}m_{i}^{2}d_{i}=1+1+1+1+2^{2}\cdot 2=12$ multiplications for $\mathrm{D}_{8}$. From a signal processing perspective, this is analogous to the computational saving of convolution being point-wise multiplications in frequency space. We visualize the computational savings obtained by working in the Fourier domain of $\mathrm{D}_{8}$ in Figure 2.
Example 3.4 (Images).
Square images can be considered as elements of $\mathbb{R}^{3\times M\times M}$ where $3$ is the number of color channels and $M$ is the image height/width in pixels. There is a natural $\mathrm{D}_{8}$-representation $\rho_{\text{image}}$ associated with square images, where $\rho_{\text{image}}(r)$ is the pixel permutation rotating the images anti-clockwise by 90° and $\rho_{\text{image}}(s)$ is the permutation reflecting the images left-to-right.
Example 3.5 (ViT features).
In ViTs, features are elements of $\mathbb{R}^{C\times N\times N}$, which we will think of as $C\times N^{2}$ matrices. Here, $N$ is the number of image tokens along the height/width of the image, so $N=M/P$ where $P$ is the patch size (typically $P=14$ or $P=16$) and $C$ is the channel dimension. The simplest representation that we consider on features is the permutation representation $\rho_{\text{token}}$ that, analogously to $\rho_{\text{image}}$, permutes the tokens according to elements of $\mathrm{D}_{8}$.
Example 3.6 (Steerable ViT features).
We can equip the channel dimension of ViT features with a group representation $\rho_{\text{chan}}$ to obtain “steerable” features. If $C$ is divisible by $8$, we can consider multiples of $\rho_{\text{reg}}$ or $\rho_{\text{iso}}$ as $\rho_{\text{chan}}$. The complete representation $\rho$ acting on the $C\times N^{2}$-matrix $\mathbf{x}$, is then permuting the tokens according to $\rho_{\text{token}}$ and modifying the channels according to $\rho_{\text{chan}}$. Concretely,
$\rho(g)\mathrm{Vec}(\mathbf{x})=\mathrm{Vec}\left(\rho_{\text{chan}}(g)\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}}\right)=\left(\rho_{\text{token}}(g)\otimes\rho_{\text{chan}}(g)\right)\mathrm{Vec}(\mathbf{x}),$ (6)
where $\mathrm{Vec}(\mathbf{x})$ is the column-wise vectorization of the matrix $\mathbf{x}$ and $\otimes$ is the tensor product of representations or (equivalently) the Kronecker product of matrices. We refer to features transforming according to equation 6 as $\rho_{\text{chan}}$-steerable, or features of type $\rho_{\text{chan}}$. This is a simpler form of the induced representations typically considered in steerable CNNs (Cohen & Welling, 2017; Weiler & Cesa, 2019), the simplification coming from the fact that we don’t enforce translation equivariance. Steerable ViT-features are illustrated in the Appendix, Figure 5.
Example 3.7 (Patchification of images).
One can consider a patchified image as a steerable ViT feature in the following way. By patchification we mean the operation of reshaping a $3\times M\times M$ image first to $N^{2}$ patches of size $P\times P$, with $NP=M$ and then to a $3P^{2}\times N^{2}$ matrix. When we transform the original image by $\rho_{\text{image}}$, the patchified image is rotated by $\rho_{\text{token}}$ and $\rho_{\text{chan}}$ as in equation 6. Now $\rho_{\text{chan}}(g)$ is a permutation matrix that rotates or mirrors a patch and we will denote this particular $\rho_{\text{chan}}$ by $\rho_{\text{patch}}$.
3.1.1 Notation
For the convenience of the reader, we collect the most important notation in the paper in this section. We use the bold letter $\mathbf{x}$ for ViT features, which have shape $C\times L$, where $L$ is the number of tokens and $C$ the channel dimension. Typically, $L=N^{2}$, where $N$ is the image height/width in tokens, or $L=N^{2}+1$ with a class token. For an individual $C$-dimensional token we use the letter $x$ which is often acted on by $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$, in which case we can split $x$ into $C/8$-dimensional sub-tokens $x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E11}},x_{\text{E12}},x_{\text{E21}},x_{\text{E22}}$, where the first four transform according to the irreps $\rho_{\text{A1}},\rho_{\text{A2}},\rho_{\text{B1}},\rho_{\text{B2}}$ respectively while the $2\times\frac{C}{8}$-matrices $(x_{\text{E11}}\quad x_{\text{E12}})^{\mathsf{T}}$ and $(x_{\text{E21}}\quad x_{\text{E22}})^{\mathsf{T}}$ both transform according to $\rho_{\text{E}}$.
3.2 Octic ViTs
We construct ViT versions that map images to steerable ViT features $\mathbf{x}\in\mathbb{R}^{C\times L}$ in the first layer (PatchEmbed) and then process steerable ViT features in subsequent layers. We choose $\rho_{\text{chan}}$ to be a multiple of $\rho_{\text{iso}}$, enabling efficient linear layers. Typically, we write $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$ where $C$ is the embedding dimension of the ViT. For classification tasks, we map the steerable ViT features to $\mathrm{D}_{8}$ invariant features fed into a classification head. We denote ViTs that use $\rho_{\text{chan}}$-steerable features in all layers as $\mathrm{D}_{8}(\text{ViT})$. We will also consider networks that use $\rho_{\text{chan}}$-steerable features for the first layers and then map either to $D\rho_{\text{A1}}$-steerable features, these networks are denoted $\mathcal{I}_{8}(\text{ViT})$, or break equivariance, denoted $\mathcal{H}_{8}(\text{ViT})$. We refer to ViTs that fall into these three families broadly as octic ViTs. In Section 4.3, we study the effect of the number of octic layers.
A commonly appreciated fact is that Transformer components $b$ such as MLP and Attention are permutation equivariant over tokens, which implies that
$b(\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}})=b(\mathbf{x})\rho_{\text{token}}(g)^{\mathsf{T}}.$ (7)
For $b$ to be fully equivariant it also needs to be equivariant in the channel dimension
$b\left(\rho_{\text{chan}}(g)\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}}\right)=\rho_{\text{chan}}(g)b(\mathbf{x})\rho_{\text{token}}(g)^{\mathsf{T}}.$ (8)
Designing the components of octic ViT blocks so that they satisfy equation 8 is the topic of Section 3.3.
Figure 3: Architecture. Patches are first extracted from an image using specialized octic filters and the resulting features are processed by $k$ octic ViT blocks. The final embeddings can be fed to lkl-k standard Transformer blocks (as demonstrated by our $\mathcal{H}_{8}$ and $\mathcal{I}_{8}$ ViTs). When $k=l$, we denote $\mathcal{I}_{8}(\text{ViT})$ by $\mathrm{D}_{8}(\text{ViT})$, which hence only uses octic ViT blocks before a final invariantization.
3.3 Octic Layers
In this section we detail our implementations of $\mathrm{D}_{8}$ equivariant Transformer layers. Together, these pieces can be combined into a ViT block $b=b_{2}\circ b_{1}$ where
$b_{1}(\mathbf{x})=\mathbf{x}+\mathrm{MHA}(\mathrm{LN}(\mathbf{x})),\quad b_{2}(\mathbf{x})=\mathbf{x}+\mathrm{MLP}(\mathrm{LN}(\mathbf{x})).$ (9)
LN is layer normalization, MHA multi-head self-attention and MLP a $C\to 4C$ linear map, GELU, and $4C\to C$ linear map. The blocks can subsequently be stacked, as illustrated in Figure 3.
3.3.1 The Patch Embedding Layer
The first layer in a ViT, following (Dosovitskiy et al., 2021), is the patch embedding, short PatchEmbed. In our case, it can be viewed as a mapping from steerable features of type $\rho_{\text{patch}}$ to steerable features of type $\frac{C}{8}\rho_{\text{iso}}$ where $C$ is the embedding dimension of the ViT. It is implemented as a convolution over the input image with kernel size and stride equal to the patch size $P$. The convolution kernels are weight sharing constrained to map to features of the different irreps-types in $\frac{C}{8}\rho_{\text{iso}}$, see Appendices C and C.1 for visualizations of the kernels.
Directly after PatchEmbed, we add a learnable positional encoding $\mathbf{e}\in\mathbb{R}^{C\times L}$ to the features. The positional encoding is not constant over tokens, thereby breaking translation equivariance. To be $\mathrm{D}_{8}$ equivariant $\mathbf{e}$ must satisfy
$\begin{split}&\rho_{\text{chan}}(g)\mathbf{x}\rho_{\text{token}}(g)^{\mathsf{T}}+\mathbf{e}=\rho_{\text{chan}}(g)(\mathbf{x}+\mathbf{e})\rho_{\text{token}}(g)^{\mathsf{T}}\quad\Longleftrightarrow\quad\mathbf{e}=\rho_{\text{chan}}(g)\mathbf{e}\rho_{\text{token}}(g)^{\mathsf{T}}.\end{split}$ (10)
[…]
3.3.2 Linear Layers
Linear layers appear in ViTs both in the MLP block and the MHA block. As mentioned in Section 3.1 and illustrated in Figure 2, equivariant linear layers map between irreps of the same type due to Schur’s lemma. This fact was used to construct efficient reflection-equivariant neural networks by Bökman et al. (2025). Here, we use the same approach for octic equivariance.
To re-iterate, for features of type $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$ we consider each token $x\in\mathbb{R}^{C}$ split into $x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E}}$ of dimensions $C/8,C/8,C/8,C/8$ and $2\times C/4$. Linear layers map each irrep type to itself, meaning that they are parameterised by four $C/8\times C/8$ matrices and one $C/4\times C/4$ matrix, yielding a factor $8$ fewer parameters than a general linear layer.
In terms of FLOPs needed to compute the linear layer, $x_{i}\mapsto W_{i}x_{i}$ requires $C/8\cdot C/8=C^{2}/64$ FLOPs for $i\in\{\text{A1, A2, B1, B2}\}$ while due to the “weight-sharing” over the two dimensions in $\rho_{\text{E}}$, it requires $2\cdot C/4\cdot C/4=C^{2}/8$ FLOPs for $i=\text{E}$. In total we therefore get $16/3\approx 5.33$ times fewer FLOPs than the $C^{2}$ required for an ordinary linear layer.
3.3.3 Activation Functions, Layer Norm, Attention and Invariantization
A pointwise activation function $\sigma$ can be applied equivariantly after transforming the features from the Fourier domain (multiples of $\rho_{\text{iso}}$), to the spatial domain (multiples of $\rho_{\text{reg}}$), as discussed in Example 3.3. In the spatial domain $\sigma$ can be applied point-wise as this commutes with the permutation representation $\rho_{\text{reg}}$. In our ViTs, following prior work, we apply the GELU activation function.
We implement (token-wise) equivariant layer normalization by transforming each irrep separately to mean 0 followed by division by the total norm over all the irreps.
If two tokens $q$ and $k$ transform according to the same orthogonal representation $\rho_{\text{chan}}$, then $q^{\mathsf{T}}k$ is invariant under $\mathrm{D}_{8}$ since $(\rho_{\text{chan}}(g)q)^{\mathsf{T}}(\rho_{\text{chan}}(g)k)=q^{\mathsf{T}}\rho_{\text{chan}}(g)^{\mathsf{T}}\rho_{\text{chan}}(g)k=q^{\mathsf{T}}k$. This means that the computation of attention logits in ordinary scaled dot-product attention is invariant, so the subsequent weighted sum over value tokens is equivariant.
To output $\mathrm{D}_{8}$ invariant classification predictions, we map from features of type $\frac{C}{8}\rho_{\text{iso}}$ to features of type $C\rho_{\text{A1}}$ and then extract the [CLS] token. We ablate different invariantizations to A1-tokens in Appendix D, including linear invariants, triple correlation (Kakarala, 2012; Sanborn & Miolane, 2023), max filtering (Cahill et al., 2024), generators of the ring of invariant polynomials and canonisation of the signal (Kaba et al., 2023). We find that a power spectrum invariantization works well and settle on that for the remainder of the experiments.
3.4 Computational Efficiency
As Transformers scale, the linear layers dominate the execution time (Kaplan et al., 2020). Thus, as ViTs grow, the FLOPs savings will approach those of the linear layer, a reduction of $5.33$ times. We plot the FLOPs savings of a ViT block as the embedding dimension increases in Figure 4(a). The computational benefits of octic ViTs are more pronounced at scale, and we benchmark the throughput of various ViT sizes from the literature and their octic counterparts in Table 1. We further compare the arithmetic intensity of standard and octic linear layers in Appendix B.
Table 1: Compute scaling. We measure the scaling of octic ViTs. The model sizes are taken from Dehghani et al. (2023) and we do not train the largest models as part of this work. The numbers ending in “x” describe the improvement over standard ViT statistics of the corresponding octic ViTs.
Model Width Depth MLP dim Heads throughput FLOPs Peak Mem.
ViT-L 1024 24 4096 16 1.32x 4.58x 2.44x
ViT-H 1280 32 5120 16 1.47x 4.58x 2.88x
ViT-G 1664 48 8192 16 1.91x 4.88x 4.36x
ViT-e 1792 56 15360 16 2.37x 5.01x 5.30x
ViT-22B 6144 36 24576 48 3.54x 5.18x 5.80x
As shown in Table 1, savings in FLOPs do not translate one-to-one to improvements in throughput (images per second) in our current implementation. However, it is still the case that the throughput is greatly improved. Our models are pure PyTorch with torch.compile, except for the GELU nonlinearity. We implement a custom Triton (Tillet et al., 2019) kernel that fuses the GELU nonlinearity with the Fourier and inverse Fourier transforms, limiting memory transfers and kernel invocation overhead. While ordinary GELU is pointwise, the new fused Triton kernel is eight to eight points. For extra efficiency, we implement the Fourier transforms by a FFT on $\mathrm{D}_{8}$, described in Appendix A.
(a) FLOPs ratio vs. embedding dimension
(b) Accuracy vs. number of octic blocks ($k$)
Figure 4: (a) Reduction in FLOPs from a non-equivariant Transformer block to an octic-equivariant block vs. embedding dimension. The matmul ratio reflects only matrix multiplications in linear layers and Attention; the total ratio includes all computations. (b) The effect of changing the number of octic blocks ($k$) for ViT-L, out of $l=24$ total blocks.
4 Experiments
In this section, we evaluate our octic ViTs on supervised (DeiT III) and self-supervised (DINOv2) training recipes and perform ablations. DeiT III (Touvron et al., 2022) is a popular and highly tuned supervised training recipe for classification and DINOv2 (Oquab et al., 2024) is a state-of-the-art self-supervised method to extract visual features at large scale. All models are trained on ImageNet-1K (Deng et al., 2009; Russakovsky et al., 2015; Recht et al., 2019) following the official implementations. We will release code and pretrained weights on GitHub.
4.1 DeiT III
We train an array of networks on the supervised task of image classification and compare to the performance reported by Touvron et al. (2022) and Bökman et al. (2025). We find, as illustrated in Table 2, that incorporating octic-equivariant layers provides significant computational savings without sacrificing accuracy. In particular, our $\mathcal{H}_{8}(\text{ViT-H/14})$ model achieves a classification performance of 85.0% compared to the baseline of 84.6% while using only 61% of the FLOPs and matching the performance of the $\mathcal{H}_{2}(\text{ViT-H/14})$ that incorporates flopping ($\mathrm{D}_{2}$) equivariance while being more computational efficient. Similar computational gains are achieved by the invariant model $\mathcal{I}_{8}(\text{ViT-H/14})$, which achieves a classification performance of 84.7%.
In the final column of Table 2, we study the effect of evaluating models on a randomly rotated validation set without training on such augmentations. We find that the invariant model performs equally well while the performance of the remaining models (including $\mathcal{H}_{8}$) significantly degrade.
Table 2: DeiT III evaluation. We measure the Top-1 classification accuracy on ImageNet-1K for different model sizes. Our networks are marked with $\dagger$.
Model params throughput FLOPs Peak Mem. Top-1 OOD Rot.
[…]
ViT-H/14 632.1 569 167.8 3285 84.6 -12.6
$\mathcal{H}_{2}(\text{ViT-H/14})$ 474.2 640 127.4 2734 85.0 -13.4
$\mathcal{I}_{8}(\text{ViT-H/14})^{\dagger}$ 362.3 657 104.0 2249 84.7 0.0
$\mathcal{H}_{8}(\text{ViT-H/14})^{\dagger}$ 355.8 660 102.3 2223 85.0 -13.4
$\mathrm{D}_{2}(\text{ViT-H/14})$ 316.1 739 87.3 2100 84.4 -13.1
$\mathrm{D}_{8}(\text{ViT-H/14})^{\dagger}$ 80.5 836 36.6 1141 82.0 0.0
ViT-L/16 304.4 1421 61.9 1557 84.2 -13.4
$\mathcal{H}_{2}(\text{ViT-L/16})$ 228.3 1610 46.9 1458 84.5 -13.9
$\mathcal{I}_{8}(\text{ViT-L/16})^{\dagger}$ 175.5 1618 38.5 1210 84.0 0.0
$\mathcal{H}_{8}(\text{ViT-L/16})^{\dagger}$ 171.3 1615 37.7 1194 84.5 -13.6
$\mathrm{D}_{2}(\text{ViT-L/16})$ 152.2 1868 32.2 1173 83.4 -14.3
$\mathrm{D}_{8}(\text{ViT-L/16})^{\dagger}$ 39.2 1877 13.5 689 79.5 0.0
Table 3: DINOv2 evaluation. We evaluate the frozen DINOv2 features by classification accuracy on ImageNet-1K (IN1K) and segmentation mIoU on ADE20K (Zhou et al., 2017b; 2019) and VOC2012 (Everingham et al., 2010). Our networks are marked with $\dagger$.
Model FLOPs IN1K (acc.) $\uparrow$ ADE20K (mIoU) $\uparrow$ VOC2012 (mIoU) $\uparrow$
($\times 10^{9}$) linear $k$-NN linear $k$-NN linear $k$-NN
ViT-H/16 127.7 81.7 81.0 34.7 30.6 70.7 60.9
$\mathcal{H}_{2}(\text{ViT-H/16})$ 97.0 81.9 80.9 34.8 30.8 70.7 61.2
$\mathcal{I}_{8}(\text{ViT-H/16})^{\dagger}$ 77.7 81.9 80.9 33.9 29.2 70.6 61.2
$\mathcal{H}_{8}(\text{ViT-H/16})^{\dagger}$ 77.5 82.2 81.4 35.1 31.1 70.8 61.7
ViT-L/16 61.9 80.9 80.5 33.2 28.4 69.1 58.1
$\mathcal{H}_{2}(\text{ViT-L/16})$ 46.9 81.3 80.7 33.4 29.0 69.3 58.0
$\mathcal{I}_{8}(\text{ViT-L/16})^{\dagger}$ 38.5 81.2 80.4 32.6 28.0 70.0 59.6
$\mathcal{H}_{8}(\text{ViT-L/16})^{\dagger}$ 37.7 81.3 80.8 33.6 29.4 69.1 59.5
$\mathrm{D}_{8}(\text{ViT-L/16})^{\dagger}$ 13.5 72.2 71.0 23.7 17.6 53.2 42.5
4.2 DINOv2
As another pretraining task, we consider the DINOv2 recipe and train our own baselines. Results are summarized in Table 3. We find that incorporating octic-equivariant layers maintains or improves downstream classification and segmentation performance while saving FLOPs. In particular, our invariant network $\mathcal{I}_{8}(\text{ViT-H/16})$ matches the downstream performance of the baseline while using only $61\%$ of the FLOPs and $\mathcal{H}_{8}(\text{ViT-H/16})$ slightly improves performance with similar savings.
In Appendix E.3, we further investigate the performance of our invariant model and evaluate it on white blood cell classification, a task that lacks a canonical orientation. The invariant model $\mathcal{I}_{8}(\text{ViT-L/16})$ outperforms the baseline on most evaluated metrics.
4.3 Ablations
Impact of equivariance.
We study the effect of equivariance by replacing the kernel constrained equivariant patch embed by an arbitrarily linear mapping while keeping the rest of the architecture the same as for $\mathcal{H}_{8}$. In principle, the model can learn to be equivariant. We evaluate ViT-B using the DeiT III recipe and achieve an accuracy of $82.4$ compared to $83.0$ for $\mathcal{H}_{8}$(ViT-B). The results suggest that equivariance yields higher accuracy than arbitrary mappings with the same block-diagonal structure (in addition to providing steerable features that can be useful in downstream applications).
Number of octic blocks.
We experiment with incorporating non-equivariant blocks (Weiler & Cesa, 2019). We include networks where the first $k$ of the ViT blocks are octic and the remaining lkl-k are standard blocks. Figure 4(b) ablates different values of $k$ for ViT-L ($l=24$). We find that $k=\frac{l}{2}$ strikes a good balance between computational efficiency and representational power and use this throughout the paper. Note that $\mathcal{I}_{8}$ performs worse for small $k$ due to early invariantization.
5 Limitations
In this work, we limit our scope to an extensive study of the $\mathrm{D}_{8}$ group and leave larger dihedral groups for future work. We follow baseline training recipes without hyperparameter tuning, and we do not conduct an extensive ablation study of the share of features per irrep or scale the size beyond ViT-H. We aim to explore this in future work. Moreover, we do not realize the full throughput potential of our octic layers, as illustrated by lower throughput gains than FLOPs savings. Our results show promise in continuing work in this direction.
6 Conclusion
We introduced octic-equivariant ViT layers that, when incorporated, maintain accuracy while significantly reducing computational complexity. We validated our proposed architectures by their effectiveness in both supervised and self-supervised learning, and conducted ablation studies to isolate the effect of invariantization, equivariance, and the number of octic blocks. In particular, we achieved an approximate 40% reduction in FLOPs for ViT-H without sacrificing accuracy, positioning octic ViTs as a strong addition to the catalog of vision architectures.
Acknowledgements
This work was supported by the Wallenberg Artificial Intelligence, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation, and by the strategic research environment ELLIIT, funded by the Swedish government. The computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at C3SE, partially funded by the Swedish Research Council through grant agreement no. 2022-06725, and by the Berzelius resource, provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre.
Appendix A Fourier Transform for $\mathrm{D}_{8}$
A.1 Implementation
The Inverse Fourier Transform, i.e., changing basis from the Isotypical representation $\rho_{\text{iso}}$ to the Regular representation $\rho_{\text{reg}}$, can be written in the case of $\mathrm{D}_{8}$ as
$Q_{\text{reg}}=\frac{\sqrt{2}}{4}\begin{pmatrix}1&1&1&1&1&1&1&-1\\ 1&1&-1&-1&1&-1&-1&-1\\ 1&1&1&1&-1&-1&-1&1\\ 1&1&-1&-1&-1&1&1&1\\ 1&-1&1&-1&-1&1&-1&-1\\ 1&-1&-1&1&-1&-1&1&-1\\ 1&-1&1&-1&1&-1&1&1\\ 1&-1&-1&1&1&1&-1&1\\ \end{pmatrix}$. (11)
In practice, we use a fast Triton-compiled implementation of the mapping $x\mapsto Q_{\text{reg}}x$ as shown in Listing LABEL:lst:q_reg, and similarly for $x\mapsto Q_{\text{reg}}^{\mathsf{T}}x$.
Listing 1: Python implementation of $Q_{\text{reg}}$.
1import math
2SQRT2_OVER_4 = math.sqrt(2) / 4
3
4def isotypical_to_regular(
5 x_A1, x_A2, x_B1, x_B2, x_E11, x_E12, x_E21, x_E22
6):
7 a = x_A1 + x_A2
8 b = x_A1 - x_A2
9 c = x_B1 + x_B2
10 d = x_B1 - x_B2
11 e = x_E11 + x_E12
12 f = x_E11 - x_E12
13 g = x_E21 + x_E22
14 h = x_E21 - x_E22
15 apc = a + c
16 amc = a - c
17 bpd = b + d
18 bmd = b - d
19 eph = e + h
20 emh = e - h
21 fpg = f + g
22 fmg = f - g
23 return (
24 SQRT2_OVER_4 * (apc + eph),
25 SQRT2_OVER_4 * (amc + fmg),
26 SQRT2_OVER_4 * (apc - eph),
27 SQRT2_OVER_4 * (amc - fmg),
28 SQRT2_OVER_4 * (bpd - fpg),
29 SQRT2_OVER_4 * (bmd - emh),
30 SQRT2_OVER_4 * (bpd + fpg),
31 SQRT2_OVER_4 * (bmd + emh)
32 )
A.2 Time complexity
[…]
Consistent with previous results, we find that the equivariant linear layers pay off more with increasing embedding dimension. The extra computations needed for the non-linearity give only a few percent overhead while the quadratic savings from the linear layers provide substantial performance gains.
[…]
Model $C=256$ $C=512$ $C=1024$ $C=2048$
GELU 44.8 82.8 155.8 302.6
[…]
MLP 128.4 324.2 1053.3 3644.9
[…]
Appendix B Arithmetic Intensity
The arithmetic intensity measures FLOPs per transferred byte and can be compared with the FLOPs per bandwidth of a given device to obtain a bound on the maximum achievable throughput (Austin et al., 2025). The arithmetic intensities are $\frac{2BCF}{P(BC+CF+BF)}$ and $\frac{2BCF/5.33}{P(BC+CF/8+BF)}$ for the standard and octic linear layers, respectively, where $B$ is the batch size in tokens, $C$ is the input dimension, $F$ is the output dimension and $P$ is the precision in bytes. This means that the octic and ordinary layers scale differently. At large scale, not only FLOPs are improved by octic layers but also arithmetic intensity. For instance, for $B=196$ (one image worth of tokens), $P=2$ and $F=4C$ (a typical MLP expansion factor) one can calculate that ordinary linear layers have higher arithmetic intensity up to $C\approx 3200$, whereas octic linear layers have higher arithmetic intensity at larger dimensions. For the experiments in this paper, we are not able to scale to such large dimensions, but still get throughput benefits due to savings in FLOPs, as shown in Table 1.
Appendix C Visualizations
We visualize the action of the octic group on images and on $\rho_{\text{iso}}$-features in Figure 5.
(a) PatchEmbed filters.
Figure 5: (a) PatchEmbed filters from a trained network. More filters are shown in Figure 6.
(b-c) Cayley diagrams showing the action of $\mathrm{D}_{8}$ on (b) patchified images and (c) $\rho_{\text{iso}}$-features. Blue arrows mean horizontal mirroring, $s$, while orange arrows mean mirroring in the bottom-left to top-right diagonal, $sr$. The features were obtained by applying the filters in (a) to the patches in (b).
(b) Octic action on patchified images.
(c) Octic action on $\rho_{\text{iso}}$-features.
C.1 Learned Filters
The PatchEmbed layer contains filters mapping the input from 3 channel dimensions to $C$ embedding dimensions. To illustrate the learned filters, we take inspiration from Dosovitskiy et al. (2021) and visualize the first 16 principal components. In contrast to regular ViTs, we have six different learned filters corresponding to the five irreps. Four for the one-dimensional irrep and two for the E irrep (due to its multiplicity). The results are illustrated in Figure 6. Interestingly, the learned filters look qualitatively different between the two learning methods. This is similar to how the learned filters of baseline DINOv2 and DeiT III look qualitatively different. DINOv2 training appears to produce more high frequency patterns while DeiT III gives clearer patterns. For the invariant irrep (A1), it appears that the DeiT III training produces a spherical pattern.
(a) Octic DeiT III Learned Filters
(a) A1
(b) A2
(c) B1
(d) B2
(e) E1
(f) E2
(b) Octic DINOv2 Learned Filters
(g) A1
(h) A2
(i) B1
(j) B2
(k) E1
(l) E2
Figure 6: Comparison of learned patch embedding filters. (a) DeiT III. (b) DINOv2. Each figure shows the top-16 principal components of the octic PatchEmbed filter for a specific feature type.
Appendix D Invariantization
There are multiple options to produce $\mathrm{D}_{8}$ invariant features, i.e. mapping tokens of type $\rho_{\text{chan}}=\frac{C}{8}\rho_{\text{iso}}$ to features of type $C\rho_{\text{A1}}$ (here denoted in short as invariantization). We let $\psi$ be a function mapping from features of type $\frac{C}{8}\rho_{\text{iso}}$ to features of type $\frac{KC}{8}\rho_{\text{A1}}$ for some $K$ which can be larger or smaller than $8$. These $KC/8$ dimensions are then mapped through a small MLP to $C$ dimensions again.
Linear Invariant (Linear).
The linear invariant simply extracts the invariant irrep. Here, $K=1$.
$\psi(x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E11}},x_{\text{E12}},x_{\text{E21}},x_{\text{E22}})=x_{\text{A1}}.$ (12)
Triple Correlation (Triple Corr.).
The triple correlation method (Sanborn & Miolane, 2023; Kakarala, 2012) extracts a complete set of third order homogeneous polynomial invariants from a signal over $\mathrm{D}_{8}$. We computed a basis for all third order invariant homogeneous polynomials using Macaulay2 (Grayson & Stillman, ; Ferraro et al., 2024) and use the basis elements as invariant. Here, $K=15$.
$\begin{split}\psi(x_{\text{A1}}&,x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E11}},x_{\text{E12}},x_{\text{E21}},x_{\text{E22}})\\ =\bigg(&x_{\text{A1}}^{3},x_{\text{A1}}(x_{\text{E21}}^{2}+x_{\text{E22}}^{2}),x_{\text{A1}}(x_{\text{E11}}x_{\text{E21}}+x_{\text{E12}}x_{\text{E22}}),x_{\text{A1}}(x_{\text{E11}}^{2}+x_{\text{E12}}^{2}),x_{\text{A1}}x_{\text{B2}}^{2},x_{\text{A1}}x_{\text{B1}}^{2},x_{\text{A1}}x_{\text{A2}}^{2},\\ &x_{\text{B2}}x_{\text{E21}}x_{\text{E22}},x_{\text{B2}}x_{\text{E12}}x_{\text{E21}}+x_{\text{B2}}x_{\text{E11}}x_{\text{E22}},x_{\text{B2}}x_{\text{E11}}x_{\text{E12}},x_{\text{B1}}x_{\text{E21}}^{2}-x_{\text{B1}}x_{\text{E22}}^{2},\\ &x_{\text{B1}}x_{\text{E11}}x_{\text{E21}}-x_{\text{B1}}x_{\text{E12}}x_{\text{E22}},x_{\text{B1}}x_{\text{E11}}^{2}-x_{\text{B1}}x_{\text{E12}}^{2},x_{\text{A2}}x_{\text{E12}}x_{\text{E21}}-x_{\text{A2}}x_{\text{E11}}x_{\text{E22}},x_{\text{A2}}x_{\text{B1}}x_{\text{B2}}\bigg)\end{split}$ (13)
Power spectrum.
A common invariant is the power spectrum. We use the following variant, with $K=6$.
$\psi(x_{\text{A1}},x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E1}},x_{\text{E2}})=\left(x_{\text{A1}},|x_{\text{A2}}|,|x_{\text{B1}}|,|x_{\text{B2}}|,\|x_{\text{E1}}\|,\|x_{\text{E2}}\|\right)$ (14)
Polynomial.
Similar to the triple correlation, we can consider a polynomial basis for the full invariant ring. This was computed using Macaulay2, yielding $K=32$.
$\begin{split}\psi(x_{\text{A1}}&,x_{\text{A2}},x_{\text{B1}},x_{\text{B2}},x_{\text{E11}},x_{\text{E12}},x_{\text{E21}},x_{\text{E22}})\\ =\bigg(&x_{\text{A1}},x_{\text{E21}}^{2}+x_{\text{E22}}^{2},x_{\text{E11}}x_{\text{E21}}+x_{\text{E12}}x_{\text{E22}},x_{\text{E11}}^{2}+x_{\text{E12}}^{2},x_{\text{B2}}^{2},x_{\text{B1}}^{2},x_{\text{A2}}^{2},\\ &x_{\text{B2}}x_{\text{E21}}x_{\text{E22}},x_{\text{B2}}x_{\text{E12}}x_{\text{E21}}+x_{\text{B2}}x_{\text{E11}}x_{\text{E22}},x_{\text{B2}}x_{\text{E11}}x_{\text{E12}},x_{\text{B1}}x_{\text{E21}}^{2}-x_{\text{B1}}x_{\text{E22}}^{2},\\ &x_{\text{B1}}x_{\text{E11}}x_{\text{E21}}-x_{\text{B1}}x_{\text{E12}}x_{\text{E22}},x_{\text{B1}}x_{\text{E11}}^{2}-x_{\text{B1}}x_{\text{E12}}^{2},x_{\text{A2}}x_{\text{E12}}x_{\text{E21}}-x_{\text{A2}}x_{\text{E11}}x_{\text{E22}},\\ &x_{\text{A2}}x_{\text{B1}}x_{\text{B2}},x_{\text{E21}}^{4}+x_{\text{E22}}^{4},x_{\text{E11}}x_{\text{E21}}^{3}+x_{\text{E12}}x_{\text{E22}}^{3},x_{\text{E11}}^{2}x_{\text{E21}}^{2}+x_{\text{E12}}^{2}x_{\text{E22}}^{2},\\ &x_{\text{E11}}^{3}x_{\text{E21}}+x_{\text{E12}}^{3}x_{\text{E22}},x_{\text{E11}}^{4}+x_{\text{E12}}^{4},\\ &x_{\text{B1}}x_{\text{B2}}x_{\text{E12}}x_{\text{E21}}-x_{\text{B1}}x_{\text{B2}}x_{\text{E11}}x_{\text{E22}},x_{\text{A2}}x_{\text{B2}}x_{\text{E21}}^{2}-x_{\text{A2}}x_{\text{B2}}x_{\text{E22}}^{2},\\ &x_{\text{A2}}x_{\text{B2}}x_{\text{E11}}x_{\text{E21}}-x_{\text{A2}}x_{\text{B2}}x_{\text{E12}}x_{\text{E22}},x_{\text{A2}}x_{\text{B2}}x_{\text{E11}}^{2}-x_{\text{A2}}x_{\text{B2}}x_{\text{E12}}^{2},\\ &x_{\text{A2}}x_{\text{B1}}x_{\text{E21}}x_{\text{E22}},x_{\text{A2}}x_{\text{B1}}x_{\text{E12}}x_{\text{E21}}+x_{\text{A2}}x_{\text{B1}}x_{\text{E11}}x_{\text{E22}},x_{\text{A2}}x_{\text{B1}}x_{\text{E11}}x_{\text{E12}},\\ &x_{\text{A2}}x_{\text{E21}}^{3}x_{\text{E22}}-x_{\text{A2}}x_{\text{E21}}x_{\text{E22}}^{3},x_{\text{A2}}x_{\text{E12}}x_{\text{E21}}^{3}-x_{\text{A2}}x_{\text{E11}}x_{\text{E22}}^{3},\\ &x_{\text{A2}}x_{\text{E11}}x_{\text{E12}}x_{\text{E21}}^{2}-x_{\text{A2}}x_{\text{E11}}x_{\text{E12}}x_{\text{E22}}^{2},x_{\text{A2}}x_{\text{E11}}^{2}x_{\text{E12}}x_{\text{E21}}-x_{\text{A2}}x_{\text{E11}}x_{\text{E12}}^{2}x_{\text{E22}},\\ &x_{\text{A2}}x_{\text{E11}}^{3}x_{\text{E12}}-x_{\text{A2}}x_{\text{E11}}x_{\text{E12}}^{3}\bigg)\end{split}$ (15)
Max filtering.
We follow Cahill et al. (2024) and implement a version of their max filtering invariant. For this, we have a set of $2C$ learnable $C$-dimensional tokens $\mathbf{y}\in\mathbb{R}^{2C\times C}$, and the $2C$ invariants are given by
$\psi(x)=\oplus_{k=1}^{2C}\max_{g\in\mathrm{D}_{8}}\langle\mathbf{y}_{k},\rho_{\text{chan}}(g)x\rangle.$ (16)
Canonisation.
Similar to the max filtering approach, we implement a canonisation where we have a single learnable $C$-dimensional reference token $y$ and compute the $C$-dimensional invariant as
$\psi(x)=\rho_{\text{chan}}\left(\mathrm{argmax}_{g\in\mathrm{D}_{8}}\langle y,\rho_{\text{chan}}(g)x\rangle\right)x.$ (17)
We conduct a study of the effect of these different invariantization methods. A priori, max filtering and canonisation should be more expressive than the others as they are the only invariants considered here that are able to preserve the relative phase information coming from different phase in different copies of $\rho_{\text{iso}}$. We train $\mathrm{D}_{8}(\text{ViT-L/16})$ on ImageNet-1K following the DeiT III recipe. The results are presented in Table 5. The conclusion is that the simple power spectrum invariant works well, and so we select it as our invariantization of choice in the remainder of the experiments.
Table 5: Invariantization Ablation. Comparing classification accuracy using different invariantization methods on ImageNet-1K using the DeiT III training recipe for 400 epochs for $\mathrm{D}_{8}$ (ViT-L/16).
Top-1 Top-5
Linear 79.2 94.7
Polynomial 79.4 94.8
Triple Corr. 76.8 93.7
Canonisation 79.5 94.9
Max Filtering 77.9 93.9
Power Spec. 79.5 94.9
Appendix E Experimental Setting
E.1 DeiT III
We train for 400 epochs on ImageNet-1K with an effective batch size of 2048 following Touvron et al. (2022) and Bökman et al. (2025). We compare to the figures reported in the respective papers and thus only train the octic ViTs. The training recipe includes heavy data augmentation (e.g. cutmix, mixup and color jitter) and uses the deprecated NVIDIA Apex library. Training is done in mixed precision with the lamb optimizer. The most important hyperparameters are summarized in Table 6(a). For more details, we refer to (Touvron et al., 2022; Bökman et al., 2025) and the official repo used for reproduction https://github.com/facebookresearch/deit.
Out of distribution (OOD) rotation evaluation simply adds a random 90 degree rotation to the validation set of IN1K and computes the classification accuracy on the randomly rotated dataset. Note, as the publicly available weights are trained for 800 epochs (whereas we compare to the figures reported for 400 epochs in the original paper), we compute the OOD $\Delta$ on these weights.
Table 6: Hyperparameters used in experiments. Collection of the most important hyperparameters used. For the full set we refer to the original implementations.
(a) DeiT III hyperparameters
Model ViT-L/16 ViT-H/14
Batch size $128\times 16$ $64\times 32$
Optimizer LAMB LAMB
LR $3\times 10^{-3}$ $3\times 10^{-3}$
LR decay cosine cosine
Weight decay 0.02 0.02
Training epochs 400 400
Warmup epochs 5 5
Stoch. Depth 0.4 0.5
Repeated Aug
Gradient Clip. 1.0 1.0
Mixup alpha 0.8 0.8
Cutmix alpha 1.0 1.0
ColorJitter 0.3 0.3
Loss BCE BCE
(b) DINOv2 hyperparameters
Model ViT-L/16 ViT-H/16
Batch size $64\times 16$ $32\times 32$
Optimizer AdamW AdamW
Base LR $4\times 10^{-3}$ $4\times 10^{-3}$
Init layer scale $10^{-5} 1.1$$10^{-5}$
Weight decay 0.04 0.04
Training steps 125K 125K
Warmup steps 12.5K 12.5K
Drop path 0.3 0.3
K 65536 65536
Gradient Clip. 3.0 3.0
Init EMA momentum 0.8 0.8
Koleo loss weight 0.1 0.1
iBOT sample prob. 0.5 0.5
iBOT mask ratio 0.1-0.5 0.1-0.5
E.2 DINOv2
We closely follow the implementation in the original paper, only modifying to train in BF16 instead of FP16 for greater stability, and follow the same evaluation protocol for classification. Lacking an official reproduction of the segmentation protocol in DINOv2, we opt for the evaluation protocol created by Darcet et al. (2025) for semantic segmentation on ADE20K and VOC2012. In contrast to the original DINOv2 paper, we decide to limit our study to ViT-L and ViT-H, the latter of which was not included in the original paper. We opt for the larger patch size of $P=16$ for all our DINOv2 models to save computational resources. Note, this is the reason why we report fewer FLOPs for our DINOv2 ViT-H models than their DeiT III counterpart (which use $P=14$).
We train on ImageNet-1K for 125K steps with an effective batch size of 1024 using the adamw optimizer. We train our own baselines for fair comparison (to obtain checkpoints after only training on IN1K). The training progression of the ViT-L/16 family can be visualized in Figure 7. The most important hyperparameters are summarized in Table 6(b). For exact details about the configuration of hyper parameters we refer to the base configs in the DINOv2 repo https://github.com/facebookresearch/dinov2.
We do not implement specific hardware efficient layers for DINOv2 training and instead opt for the standard octic layers that are timm compatible. As such, the octic layers do not leverage NestedTensorBlock and training speedups associated with xFormers. This choice does not impact speed on downstream tasks but slightly decreases pre-training speed.
(a) linear
(b) $k$-NN
Figure 7: DINOv2 training progression. Classification accuracy development during 125K training steps for linear probe and $k$-NN on frozen features for ViT-L sized models.
We extend the evaluation of DINOv2 (trained on ImageNet-1K) to two popular classification datasets and report the performance in Table 7. We find, once again, that our networks achieve similar or better performance than the baseline while using substantially fewer FLOPs.
Table 7: DINOv2 additional evaluation. We further evaluate the frozen DINOv2 features by classification accuracy on iNaturalist2021 (Van Horn et al., 2021) and Places365 (Zhou et al., 2017a).
Model FLOPs iNaturalist2021 $\uparrow$ Places365 $\uparrow$
($\times 10^{9}$) linear $k$-NN linear $k$-NN
ViT-H/16 127.7 81.7 81.0 34.7 30.6
$\mathcal{I}_{8}(\text{ViT-H/16})$ 77.7 81.9 80.9 33.9 29.2
$\mathcal{H}_{8}(\text{ViT-H/16})$ 77.5 82.2 81.4 35.1 31.1
ViT-L/16 61.9 80.9 80.5 33.2 28.4
$\mathcal{I}_{8}(\text{ViT-L/16})$ 38.5 81.2 80.4 32.6 28.0
$\mathcal{H}_{8}(\text{ViT-L/16})$ 37.7 81.3 80.8 33.6 29.4
E.3 DinoBloom
We investigate the performance of our invariant model on white blood cell classification. We follow the procedure of DinoBloom (Koch et al., 2024) and report our results for ViT-L in Table 9. We find that the invariant model $\mathcal{I}_{8}(\text{ViT-L/16})$ outperforms the baseline on most evaluated metrics. We tried evaluating on a rotated test set and found negligible change in performance for the baseline (the invariant model inherently has no change in performance, similar to the last column of Table 2).
[…]
Table 8: Hematology datasets. Dataset mixture for DinoBloom finetuning.
Datasets Modality Images
BMC Bone Marrow 171K
AML Hehr Blood 102K
APL Blood 26K
AML Matek Blood 18K
Acevedo Blood 17K
Raabing WBC Blood 10K
Total 354k
Table 9: Hematology finetuning. White blood cell classification performance on BMC dataset with 21 highly imbalanced classes after finetuning on hematology data following DinoBloom.
Model 1-$k$-NN 20-$k$-NN Lin. probe
wF1 Acc bAcc wF1 Acc bAcc wF1 Acc bAcc
ViT-L/16 78.0 78.0 57.6 83.1 83.6 54.9 84.6 84.7 62.2
$\mathcal{I}_{8}(\text{ViT-L/16})$ 78.1 78.0 61.3 83.5 83.9 55.7 85.0 85.2 61.9
E.4 General Settings
Software versioning.
We utilize the PyTorch (Paszke et al., 2019) and the timm (Wightman, 2019) libraries for our experiments. We run the same versioning as our benchmarks. For all other experiments, we use Python 3.11.9 and PyTorch 2.6.0 with CUDA 11.8.
Model sizes.
The model sizes referred to in the paper adhere to the standard terminology used by Wightman (2019); Dosovitskiy et al. (2021). If we denote the shape by a tuple of (depth, width, attention heads), ViT-L has shape (24, 1024, 16) and ViT-H has shape (32, 1280, 16). Both use MLP dimension four times the size of the embedding dimension (commonly referred to as MLP ratio).
Calculating throughput.
Throughput and peak memory are measured on a single A100-80GB GPU with batch size fixed to 64 using torch.compile, FlashAttention (Dao, 2024), and mixed precision. The throughput only measures forward passes with no gradients. Moreover, we utilize 10 warm-up iterations and then average over 100 runs (Bökman et al., 2025). Peak memory is measured with PyTorch’s device memory allocation monitor.
Counting FLOPs.
We count the number of FLOPs using fvcore.nn.FlopCountAnalysis https://github.com/facebookresearch/fvcore. FLOPs are normalized with respect to the batch size (i.e. we measure FLOPs/image). We acknowledge that the term FLOPs often leads to confusion. We adopt the terminology of prior work (Touvron et al., 2022; Bökman et al., 2025) and the fvcore library for FLOPs, though, strictly speaking, this refers to MACs (as a factor of two is omitted).
E.5 Compute Resources
Table 10 provides information on the computing resources required for the main experiments. The largest single experiment (i.e. not accounting ablation studies) used $32\timesA$100-40GB for 58 hours. Failed or discarded results are not included. In total, we used around 20k A100-40GB hours for our main results.
Table 10: GPU hours. Exact account of hardware usage for the main experiments. Account does not include failed or discarded experiments or smaller evaluation pipelines such as linear probes and $k$-NN. Efficiency gains were made progressively and thus some experiments took longer than necessary. Similarly, our training pipeline for DINOv2 is not fully optimized for the octic layers. Resource unit (RU) is measured in A100 equivalent hours where an A40 hour costs 0.54 units.
Experiment Model GPUs Time RU
DeiT III $\mathcal{I}_{8}$(ViT-H/14) $32\timesA$100-40GB 58h 1856
DeiT III $\mathrm{D}_{8}$(ViT-H/14) $32\timesA$100-40GB 57h 1824
DeiT III $\mathcal{H}_{8}$(ViT-H/14) $32\timesA$100-40GB 55h 1760
DeiT III $\mathcal{H}_{8}$(ViT-L/16) $16\timesA$100-40GB 39h 624
DeiT III $\mathcal{I}_{8}$(ViT-L/16) $16\timesA$100-40GB 37h 592
DeiT III $\mathrm{D}_{8}$(ViT-L/16) $16\timesA$100-40GB 36h 576
Ablation: Invarisation $\mathcal{H}_{8}$(ViT-L/16) $16\timesA$100-40GB 240h 3840
Ablation: Hybridisation $\mathcal{I}_{8}$(ViT-L/16) $16\timesA$100-40GB 122h 1952
Ablation: Hybridisation $\mathcal{H}_{8}$(ViT-L/16) $16\timesA$100-40GB 120h 1920
DINOv2 $\mathcal{H}_{8}$(ViT-H/16) $32\timesA$100-40GB 33h 1056
DINOv2 $\mathcal{I}_{8}$(ViT-H/16) $32\timesA$100-40GB 30h 960
DINOv2 $\mathrm{D}_{8}$ (ViT-L/16) $32\timesA$100-40GB 23h 736
DINOv2 ViT-H/16 $32\timesA$100-40GB 21h 672
DINOv2 $\mathcal{I}_{8}$(ViT-L/16) $32\timesA$100-40GB 19h 608
DINOv2 $\mathcal{H}_{8}$(ViT-L/16) $32\timesA$100-40GB 19h 608
DINOv2 ViT-L/16 $16\timesA$100-40GB 24h 384
DinoBloom $\mathcal{I}_{8}$(ViT-L/16) $8\timesA$100-40GB 46min 6
DinoBloom ViT-L/16 $8\timesA$100-40GB 37min 5
Total 19979
Appendix F Licenses
Data.
All our models are trained on ImageNet-1K (Deng et al., 2009) which has a custom non-commercial license (see https://www.image-net.org/download.php). We also evaluate ADE20K, VOC2012, and BMC. For the licenses, see https://ade20k.csail.mit.edu/terms/index.html, http://host.robots.ox.ac.uk/pascal/VOC/voc2012/, and https://www.cancerimagingarchive.net/data-usage-policies-and-restrictions/, respectively.
Images.
All the images in this paper are original and taken by the authors. Similarly, illustrations are created by the authors. The assets allow for non-commercial use and redistribution with proper attribution (CC BY-NC).
Code.
The efficient octic layers are our original work and will be licensed under Apache License 2.0 following the code release. The training pipelines are from DeiT III (Touvron et al., 2022) and DINOv2 (Oquab et al., 2024), which are also under Apache License 2.0. For details about the Apache License 2.0 we refer to https://www.apache.org/licenses/LICENSE-2.0.